OpenStudy (anonymous):

last question i promise !! solve the inequality 5x - 3 <= 2x + 10 describe the graph ?

OpenStudy (anonymous):

ok lets do it 5x-3 <= 2x+10 5x-2x <=10+3 3x <= 13 x <= 13/3 ---->solved

OpenStudy (anonymous):